An amount of R3 450 earns simple interest and accumulates to R5623, 50 after seven years. Had the yearly interest rate been 2% more, how much interest would he have accumulated in seven years


1
Expert's answer
2021-05-17T16:07:20-0400

"5623.50 -3450 = 3450 \times r_1 \times 7 \\

2173.5 = 24150 \times r_1 \\

r_1 = \frac{2173.5}{24150} = 0.09 = 9 \; \% \\

r_2 = r_1 + 2 = 11 \; \% = 0.11 \\

X-3450 = 3450 \times 0.11 \times 7 \\

X-3450 = 2656.5 \\

X = 6106.5 \\

Interest = 6106.5 -3450 = 2656.5"